# Detect, Investigate, Judge and Determine: A Novel LLM-based Framework for Few-shot Fake News Detection

### v1

@@ Abstract
Few-Shot Fake News Detection (FS-FND) aims to distinguish inaccurate news from real ones in extremely low-resource scenarios. This task has garnered increased attention due to the widespread dissemination and harmful impact of fake news on social media. Large Language Models (LLMs) have demonstrated competitive performance with the help of their rich prior knowledge and excellent in-context learning abilities. However, existing methods face significant limitations, such as the Understanding Ambiguity and Information Scarcity, which significantly undermine the potential of LLMs. To address these shortcomings, we propose a Dual-perspective Augmented Fake News Detection (DAFND) model, designed to enhance LLMs from both inside and outside perspectives. Specifically, DAFND first identifies the keywords of each news article through a Detection Module. Subsequently, DAFND creatively designs an Investigation Module to retrieve inside and outside valuable information concerning to the current news, followed by another Judge Module to derive its respective two prediction results. Finally, a Determination Module further integrates these two predictions and derives the final result. Extensive experiments on two publicly available datasets show the efficacy of our proposed method, particularly in low-resource settings.
@@ 1 Introduction
Figure 1: An example of fake news detection and the comparison between existing LLM-based methods and our design.
@@
Generally, fake news detection can be framed as a binary classification problem and addressed using various classification models. In the early stage, researchers primarily employ machine learning or deep learning algorithms to represent and classify candidate news articles (Horne and Adali, 2017; Jiang et al., 2022). More recently, with the advent of Large Language Models (LLMs), FS-FND has been effectively addressed through the In-Context Learning (ICL) technology, which is particularly prevalent in few-shot settings (Hu et al., 2024; Boissonneault and Hensen, 2024). Among them, Hu et al. (2024); Wang et al. (2024); Teo et al. (2024) were pioneers in investigating the potential of LLMs in this field.
@@
However, this kind of methods mainly directly ask the LLM to judge the authenticity of the given news, which often exceeds the capabilities of LLMs in many circumstances, particularly those relatively small LLMs in common usage (e.g., 7B parameters). An example of this can be observed in Figure 1, which presents a news about microchip developments. Existing LLM-based approaches encounter two principal challenges: (1) Understanding Ambiguity: LLMs may fail to understand and grasp the core meaning conveyed in the news, thereby straining the detection process. (2) Information Scarcity: Due to the timeliness nature of news content, the training corpus of LLMs is frequently outdated. This poses fundamental difficulties in the detection of fake news.
@@
To this end, this paper proposes a novel approach to address the two aforementioned issues. Specifically, to mitigate the Understanding Ambiguity problem, we aim to extract valuable insights from an inside perspective by retrieving similar samples from the training set, thereby enhancing the comprehension of key concepts in the target news. Concurrently, to tackle the Information Scarcity problem, we employ an external search engine to gather relevant information about the news online. This integrates real-time data, effectively overcoming the limitation of information obsolescence.
@@
More specifically, we design a Dual-perspective Augmented Fake News Detection (DAFND) model. DAFND comprises four key components: (a) A Detection Module: This module employs LLMs to extract keywords from each news article, providing an effective query for the subsequent modules. (b) An Investigation Module: It investigates more valuable information related to the target news, which comes from both inside (i.e., training set) and outside (i.e., search engine) perspectives. (c) A Judge Module: This module designs prompts that enable LLMs to generate predictions and explanations based on the findings from both inside and outside investigations. (d) A Determination Module: It takes into account the predictions and explanations from both two perspectives, and thus makes a final decision with high confidence, especially in cases where two perspectives conflict.
In summary, the main contributions of our work could be summarized as follows.
  • For the first time, we explore augmenting LLMs from inside and outside perspectives for few-shot fake news detection, pioneering a novel direction in this field.
  • We devise the Dual-perspective Augmented Fake News Detection (DAFND) model, which effectively addresses the Understanding Ambiguity and Information Scarcity problems, particularly in low-resource settings.
@@
  • We conduct extensive experiments on two publicly available datasets, where the experimental results demonstrate the effectiveness of our proposed method. We will make our codes publicly available upon acceptance of the paper.
@@ 2.1 Few-Shot Fake News Detection
The objective of fake news detection task is to distinguish inaccurate news from real ones (Shu et al., 2017). For few-shot fake news detection, only $K$ instances per category ($K$-shot) are sampled for the training phase (Gao et al., 2021; Ma et al., 2023).
@@
Generally, fake news detection can be defined as a binary classification problem and is addressed by a variety of classification models. Initially, researchers mainly rely on feature engineering and machine learning algorithms. For example, Horne and Adali (2017) presented a set of content-based features to a Support Vector Machine (SVM) classifier. With the rapid development of computing power, significant improvements have been made with the help of various deep learning algorithms and Pre-trained Language Models (PLMs). Ghanem et al. (2021) combined lexical features and a Bi-GRU network to achieve accurate fake news detection. Jiang et al. (2022) proposed Knowledgeable Prompt Learning (KPL), incorporating prompt learning into fake news detection for the first time, and achieved state-of-the-art performance.
@@
Additionally, due to the specificity of news articles, researchers have incorporated external knowledge knowledge to contribute to traditional fake news detection models. For instance, Dun et al. (2021); Ma et al. (2023); Hu et al. (2021b) utilized knowledge graphs to enrich entity information and structured relation knowledge, leading to more precise news representations and improved detection performance. Meanwhile, Huang et al. (2023) adopted a data augmentation perspective, proposing a novel framework for generating more valuable training examples, which has proven to be beneficial in detecting human-written fake news.
@@
Currently, with the advent of large language models, many researchers are exploring few-shot fake news detection through in-context learning and data augmentation technologies (Hu et al., 2024; Wang et al., 2024; Teo et al., 2024). However, These methodologies either simply apply LLMs to judge the authenticity of the given news, or employ LLMs to rephrase the training data, thereby not fully utilizing the potential of LLMs. More importantly, most of them are significantly limited by the aforementioned two shortcomings, particularly in the Information Scarcity problem.
Figure 2: The architecture of our DAFND model. It includes four sequentially connected parts: (a) Detection Module; (b) Investigation Module; (c) Judge Module; (d) Determination Module.
2.2 Large Language Models
@@
The emergence of Large Language Models (LLMs) such as GPT-4, LLama-3 and others (Hoffmann et al., 2022; OpenAI, 2023; AI@Meta, 2024; Tunstall et al., 2023), marks a significant advancement in the field of natural language processing. In-context learning, a novel few-shot learning paradigm, was initially introduced by Brown et al. (2020). To date, LLMs have exhibited remarkable performance across a range of NLP tasks, including text classification, information extraction, question answering and fake news detection (Hu et al., 2024; Wang et al., 2024; Teo et al., 2024; Liu et al., 2022; Zhao et al., 2021).
Previous studies (Hu et al., 2024; Wang et al., 2024; Teo et al., 2024) have endeavored to solve few-shot fake news detection via directly asking LLMs or employing them to rephrase the training data. For example, Hu et al. (2024) explored the potential of LLMs in fake news detection, and further developed an Adaptive Rationale Guidance (ARG) network to synergize traditional methods with large language models. Similarly, Wang et al. (2024) leveraged LLMs to generate justifications towards evidence relevant to given news, which were subsequently fed into a trainable classifier.
@@ 3 Problem Statement
4 The DAFND Model
In this section, we will introduce the technical details of DAFND model. As depicted in Figure 2, DAFND comprises of four distinct components: (a) Detection Module; (b) Investigation Module; (c) Judge Module; (d) Determination Module. These modules are sequentially interconnected to achieve the final accurate detection of fake news.
@@ 4.1 Detection Module
In this module, we aim to identify the key information contained in the given news article, which will serve as the query for the subsequent modules. Specifically, we construct prompts to convey the original news article to the LLM. The LLM is then guided to extract $N$ keywords $\{w_{1},w_{2},...,w_{N}\}$, which are expected to address the question: “when, where, who, what, how and why did the given news $S$ happen?”:
$\{w_{1},...,w_{N}\}=\mathcal{F}(Ins_{w},S),$ (1)
where $\mathcal{F}$ represents the LLM, and $Ins_{w}$ denotes the prompt for the in-context learning. We provide a more detailed description about it in Appendix A.1.
@@ 4.2 Investigation Module
In this module, we aim to investigate the relevant information to assist the LLM in conducting accurate inferences. As outlined in Section 1, this process is approached from two perspectives: Inside Investigation and Outside Investigation.
Inside Investigation.
To address the Understanding Ambiguity problem introduced in the Introduction, we retrieve effective demonstrations to enhance the LLM’s understanding during the in-context learning process (Liu et al., 2022; Rubin et al., 2022).
Specifically, we first concatenate the extracted $N$ keywords of each news, and then utilize the pre-trained language model $\mathcal{M}$ to obtain the representation of these keywords $\{w_{1},...,w_{N}\}$:
$W=w_{1}$ $\uplus w_{2}\uplus...\uplus w_{N},$ (2)
$H$ $=\mathcal{M}(W),$
@@
Subsequently, when inferring a candidate news $j$, we employ the $k$-Nearest Neighbors ($k$NN) search method (Khandelwal et al., 2019) to retrieve valuable samples from the training set. In detail, we use the representation $H_{j}$ of news $j$ to query the datastore $D$ according to the euclidean distance. Then, based on the computed distance, we select the nearest $k$ positive and negative news samples, respectively [^3] This approach ensures the diversity and effectiveness of the retrieved samples.:
@@
Outside Investigation.
In response to the Information Scarcity problem, we further retrieve additional real-time information from external sources. Inspired by Yoran et al. (2023); Paranjape et al. (2023), we implement a retriever based on the google search engine, using the SerpAPI service [^4] https://serpapi.com/.
Specifically, based on the extracted keywords $\{w_{1},w_{2},...,w_{N}\}$ in Section 4.1, we first concatenate them to construct the initial query $\mathcal{Q}=w_{1}\uplus w_{2}\uplus...\uplus w_{N}$. Then, following the strategy proposed in (Yoran et al., 2023), we further format the search queries as “en.wikipedia.org $\mathcal{Q}$”, with the Wikipedia domain preceding the intermediate question. We return the top-1 evidence retrieved by Google. And all retrieved evidence sentences are prepended to the outside investigation outcome, denoted as $\mathcal{OI}$.
@@ 4.3 Judge Module
Following the paradigm designed in the Investigation Module (Section 4.2), we attempt to derive the inference results based on the investigated information from both inside and outside perspectives.
Inside Judge.
After obtaining the effective demonstrations $\mathcal{II}$ from Inside Investigation, we design prompts to provide the essential information to the LLM, thereby generating the inside prediction. Specifically, inspired by the various attempts about in-context learning Paranjape et al. (2023), we first describe the target of the fake news detection task through an inside instruction. Then, the retrieved inside investigation results =$\mathcal{II}=\{U_{positive},U_{negative}\}$ of current candidate news are followed, which augment the LLM’s understanding of this task. Finally, we prompt the LLM to predict the result of current news and give its corresponding supportive explanation.
In summary, the inside judge process can be expressed as:
$P_{i},R_{i}=\mathcal{F}(Ins_{i}\ \uplus\ \mathcal{II}\ \uplus\ x_{test}),$ (4)
[…]
Outside Judge.
With the outside investigation information $\mathcal{OI}$ retrieved through the google search engine as described in Section 4.2, we can derive the outside judge prediction, which is crucial for real-time news detection.
Similar to the design of Inside Judge, we describe the objective of fake news detection through an outside instruction, followed by the candidate news to be detected and the retrieved outside investigation documents $\mathcal{OI}$. After that, we can derive the outside prediction $P_{o}$ and explanation $R_{o}$:
$P_{o},R_{o}=\mathcal{F}(Ins_{o}\ \uplus\ x_{test}\ \uplus\ \mathcal{OI}),$ (5)
where $Ins_{o}$ denotes the outside instruction. Detailed information about this prompt is available in Appendix A.3.
Following these two parallel judge processes, we obtain the prediction results $P_{i}$, $P_{o}$ and corresponding explanations $R_{i}$, $R_{o}$ from inside and outside perspectives, respectively.

### v2

# Detect, Investigate, Judge and Determine: A Knowledge-guided Framework for Few-shot Fake News Detection

@@ Abstract
Few-Shot Fake News Detection (FS-FND) aims to distinguish inaccurate news from real ones in extremely low-resource scenarios. This task has garnered increased attention due to the widespread dissemination and harmful impact of fake news on social media. Large Language Models (LLMs) have demonstrated competitive performance with the help of their rich prior knowledge and excellent in-context learning abilities. However, existing methods face significant limitations, such as the Understanding Ambiguity and Information Scarcity, which significantly undermine the potential of LLMs. To address these shortcomings, we propose a Dual-perspective Knowledge-guided Fake News Detection (DKFND) model, designed to enhance LLMs from both inside and outside perspectives. Specifically, DKFND first identifies the knowledge concepts of each news article through a Detection Module. Subsequently, DKFND creatively designs an Investigation Module to retrieve inside and outside valuable information concerning to the current news, followed by another Judge Module to evaluate the relevance and confidence of them. Finally, a Determination Module further derives two respective predictions and obtain the final result. Extensive experiments on two public datasets show the efficacy of our proposed method, particularly in low-resource settings.
Detect, Investigate, Judge and Determine:
A Knowledge-guided Framework for Few-shot Fake News Detection
Ye Liu1,2, Jiajun Zhu1,2, Xukai Liu1,2, Haoyu Tang1,2 Yanghai Zhang1,2, Kai Zhang1,2, Xiaofang Zhou3, Enhong Chen2,3 1 University of Science and Technology of China 2 State Key Laboratory of Cognitive Intelligence 3 The Hong Kong University of Science and Technology liuyer@mail.ustc.edu.cn
@@ 1 Introduction
Figure 1: An example of fake news detection and the limitations of existing LLM-based methods.
@@
Generally, fake news detection can be framed as a binary classification problem and addressed using various classification models. In the early stage, researchers primarily employ machine learning or deep learning algorithms to represent and classify candidate news articles (Horne and Adali, 2017; Jiang et al., 2022). More recently, with the rise of Large Language Models (LLMs), FSFND has been effectively addressed through the in-context learning technology, which is particularly prevalent in few-shot settings (Hu et al., 2024; Boissonneault and Hensen, 2024). Among them, Hu et al. (2024); Wang et al. (2024); Teo et al. (2024) were pioneers in investigating the potential of LLMs in this field.
@@
However, despite these advancements, existing LLM-based methods face inherent challenges. They often rely on directly prompting LLMs to judge the authenticity of the given news, which often exceed the capabilities of these models, particularly those relatively small LLMs (e.g., 7B parameters) that are commonly used in practical applications. An example of this can be observed in Figure 1, which presents a news about microchip developments. Existing LLM-based approaches encounter two principal challenges: (1) Understanding Ambiguity: LLMs may fail to understand and grasp the core meaning conveyed in the news, especially when it contains subtle nuances or domain-specific jargon, thereby straining the detection process. (2) Information Scarcity: Given the dynamic and fast-paced nature of news, the training corpus of LLMs is often outdated, leading to a lack of relevant and up-to-date information during the detection process.
@@
To this end, this paper proposes a novel approach to address the two aforementioned issues. Specifically, to mitigate the Understanding Ambiguity problem, we aim to extract valuable insights from an inside perspective by retrieving similar samples from the training set, thereby enhancing the comprehension of key concepts in the target news. Simultaneously, to tackle the Information Scarcity problem, we employ an external search engine to gather relevant information about the news from the web. This design integrates more real-time data, effectively overcoming the limitation of information obsolescence. Importantly, in the whole process, an external knowledge graph is employed to enhance the reliability and information relevance.
@@
More specifically, we design a Dual-perspective Knowledge-guided Fake News Detection (DKFND) model. DKFND comprises four key components: (a) A Detection Module: This module leverages a knowledge graph to identify relevant knowledge concepts within the text, forming the basis for queries used in the subsequent modules. (b) An Investigation Module: It investigates more valuable information related to the target news, which comes from both inside (i.e., training set) and outside (i.e., search engine) perspectives. (c) A Judge Module: With the help of knowledge graphs and LLMs, this module assesses the relevance and authenticity of investigated informations. (d) A Determination Module: It designs specialized prompts to guide LLMs in generating predictions and explanations from both two perspectives. Subsequently, it makes a final decision with high confidence, especially in cases where the inside and outside perspectives conflict.
@@
We conduct extensive experiments on two publicly available datasets, where the experimental results demonstrate the superiority of our proposed method. Our codes is avaiable via https://anonymous.4open.science/r/DKFND-ED55.
@@ 2.1 Few-Shot Fake News Detection
Generally, fake news detection can be defined as a binary classification problem and addressed by a variety of classification models. Initially, researchers mainly rely on feature engineering and machine learning algorithms (Horne and Adali, 2017). As computing power and data availability have increased, significant improvements have been made with the help of various deep learning algorithms and Pre-trained Language Models (PLMs). For instance, Ghanem et al. (2021) combined lexical features and a Bi-GRU network to achieve accurate fake news detection. Jiang et al. (2022) introduced the Knowledgeable Prompt Learning (KPL), a novel framework that integrates prompt learning with fake news detection, achieving state-of-the-art performance in few-shot settings.
@@
Additionally, researchers have also recognized the importance of external knowledge to complement traditional fake news detection methods. News articles often contain references to entities, events, and facts that are external to the article itself, which makes external knowledge integration crucial for accurate detection. For instance, Dun et al. (2021); Ma et al. (2023); Hu et al. (2021b) utilized knowledge graphs to enrich entity information and structured relation knowledge, leading to more precise news representations and improved detection performance. Meanwhile, Huang et al. (2023) adopted a data augmentation perspective, proposing a novel framework for generating more valuable training examples, which has proven to be beneficial in detecting human-written fake news.
@@
More recently, with the advent of large language models, many researchers are exploring few-shot fake news detection through in-context learning and data augmentation technologies (Hu et al., 2024; Wang et al., 2024; Teo et al., 2024). For example, Hu et al. (2024) explored the role of LLMs in fake news detection, proposing an Adaptive Rationale Guidance (ARG) network that combines traditional detection techniques with the generative capabilities of LLMs. However, many of these approaches rely on directly prompting the LLM to classify the news as fake or real, without fully utilizing the broader potential of LLMs to understand and reason about the news context. More importantly, most of them are significantly limited by the aforementioned two shortcomings, particularly in the Information Scarcity problem.
2.2 Large Language Models & Retrieval-Augmented Generation
@@
Large Language Models (LLMs) such as GPT-4, LLama-3, and others have revolutionized natural language processing, achieving impressive results across a variety of tasks, including text classification, summarization, information extraction, and fake news detection (Hoffmann et al., 2022; OpenAI, 2023; AI@Meta, 2024; Tunstall et al., 2023). One of the key innovations introduced by LLMs is in-context learning, a paradigm in which the model learns from a few examples presented in the context of the task, allowing it to perform few-shot learning without the need for explicit retraining (Brown et al., 2020; Hu et al., 2024; Wang et al., 2024).
In parallel, the integration of Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance the capabilities of LLMs. RAG models, such as those proposed by Lewis et al. (2020), combine the generative power of LLMs with the retrieval of relevant documents or knowledge from an external corpus. This approach not only improves the factual accuracy of generated content but also allows for more informed and contextually appropriate responses in tasks like question answering and knowledge-intensive applications (Izacard and Grave, 2021; Borgeaud et al., 2022; Izacard et al., 2023).
In the context of few-shot fake news detection, incorporating external knowledge retrieval has shown promising results. Jiang et al. (2022) were among the first to integrate retrieval-based methods with LLMs for fake news detection. By incorporating knowledge from Wikidata into the prompt representation process, they devised a knowledgeable prompt learning framework that significantly improved the detection performance. Similarly, other studies have optimized external retrieval mechanisms to enhance the performance of LLMs, allowing more context-aware and accurate predictions (Borgeaud et al., 2022; Izacard et al., 2023).
@@ 3 Problem Statement
4 The DKFND Model
In this section, we will introduce the technical details of DKFND model, as illustrated in Figure 2.
Figure 2: The architecture of our DKFND model. It includes four sequentially connected parts: (a) Detection Module; (b) Investigation Module; (c) Judge Module; (d) Determination Module.
@@ 4.1 Detection Module
In this module, we aim to identify the key information contained in the given news article, which will serve as the query for the subsequent modules. Specifically, we employ the Spacy algorithm Honnibal and Montani (2017) to extract relevant knowledge concepts $C=\{c_{1},c_{2},...,c_{m}\}$.
Subsequently, we construct prompts to guide the LLM to select the top-$N$ most critical knowledge concepts, which are expected to answer the question: “How relevant these knowledge concepts are to the given news?”:
$C^{\prime}=\mathcal{F}(P_{detection},S,C),$ (1)
where $\mathcal{F}$ represents the LLM, and $P_{detection}$ denotes the prompt for the in-context learning. More detailed description can be found in Appendix C.1.
@@ 4.2 Investigation Module
In this module, we aim to investigate the relevant information, from two perspectives: Inside Investigation and Outside Investigation.
4.2.1 Inside Investigation.
To address the Understanding Ambiguity problem introduced in Section 1, we retrieve effective demonstrations to enhance the LLM’s understanding during the in-context learning process (Liu et al., 2022; Rubin et al., 2022).
Specifically, we first concatenate the extracted $N$ concepts of each news, and then utilize the pre-trained language model $\mathcal{M}$ to obtain the representation of these concepts $\{c_{1},...,c_{N}\}$:
$Q=c_{1}$ $\uplus c_{2}\uplus...\uplus c_{N},$ (2)
$H$ $=\mathcal{M}(Q),$
@@
Subsequently, when inferring a candidate news $j$, we employ the $k$-Nearest Neighbors ($k$NN) search method (Khandelwal et al., 2019) to retrieve valuable samples from the training set. In detail, we use the representation $H_{j}$ of news $j$ to query the datastore $D$ according to the euclidean distance. Then, based on the computed distance, we select the nearest $k$ positive and negative news samples, respectively:
@@
4.2.2 Outside Investigation.
In response to the Information Scarcity problem, we further retrieve additional real-time information from external sources. Inspired by Yoran et al. (2023); Paranjape et al. (2023), we implement a retriever based on the google search engine, using the SerpAPI service [^2] https://serpapi.com/. You can refer to Section B for a searching example..
Specifically, based on the extracted concetps $\{c_{1},c_{2},...,c_{N}\}$ in Section 4.1, we first concatenate them to construct the initial query $\mathcal{Q}=c_{1}\uplus c_{2}\uplus...\uplus c_{N}$. Then, following the strategy proposed in (Yoran et al., 2023), we further format the search queries as “en.wikipedia.org $\mathcal{Q}$”, with the Wikipedia domain preceding the intermediate question. We return the top-L evidence retrieved by Google. And all retrieved evidence sentences are prepended to the outside investigation outcome, denoted as $\mathcal{OI}=\{G_{k},k=1,2,..,L\}$.
@@ 4.3 Judge Module
We design this module to assess the relevance and authenticity of the retrieved information.
4.3.1 Inside Judge
In the Inside Investigation module (Section 4.2.1), we employ the kNN search method to retrieve relevant documents $\mathcal{II}$ from the training set. However, the retrieved content may still contain irrelevant information, potentially introducing noise and interfering with the prediction process.
To address this issue, we propose an inside judge method that leverages the knowledge concepts identified in the Detection Module. Specifically, we design prompts to guide the LLM in further refining the retrieved documents by selecting the top- $m$ most relevant samples. Specifically:
$U_{positive}^{\prime}$ $=\mathcal{F}\ (P_{IJ}\ \uplus\ U_{positive}\ \uplus\ C^{\prime}),$ (4)
$U_{negative}^{\prime}$ $=\mathcal{F}\ (P_{IJ}\ \uplus\ U_{negative}\ \uplus\ C^{\prime}),$
where $P_{IJ}$ denotes the prompt and more detailed can be found in Appendix C.2. Finally, the selected inside investigated information can be denoted as =$\mathcal{II}^{\prime}=\{U_{positive}^{\prime},U_{negative}^{\prime}\}$.
4.3.2 Outside Judge
For the Ouside Investigation module (Section 4.2.2), although the retrieved information is obtained via the Google API, ensuring the relevance, a major limitation lies in the uncertainty of its authenticity and accuracy. To address this, we incorporate a knowledge graph to verify the authenticity of the retrieved outside information.
Concept Triple Identification.
In this part, we aim to identify the knowledge concept triples from each document $G_{k}$ in $\mathcal{OI}$. Specifically, we first employ Spacy algorithm to recognize the knowledge concept set $\mathcal{E}_{k}=\{e_{i}\ |\ i=1,2...\}$ from $G_{k}$. Then, we design prompts to guide the LLM to compose triples:
$\mathcal{T}_{k}=\mathcal{F}\ (P_{OJ}\ \uplus\ G_{k}\ \uplus\ \mathcal{E}_{k}),$ (5)
where $P_{OJ}$ denotes the prompt, and $\mathcal{T}_{k}=\{T_{k}^{l}|T_{k}^{l}=(h_{k}^{l},r_{k}^{l},t_{k}^{l}),l=1,2...\}$. Subsequently, to refine the relation $r_{k}^{l}$, we utilize the semantic matching method to find a closest relation in KGs. For convenience, we still use $\mathcal{T}_{k}$ to denote the refined concept triples.
Triple Authenticity Assessment.
In this part, we further employ the KG embedding model: ComplEx (Trouillon et al., 2016), which is pretrained on Wikidata through LibKGE (Broscheit et al., 2020), to assess the authenticity of each triple $T_{k}^{l}$ in $\mathcal{T}_{k}$:
$Sc(T_{k}^{l})=f(h_{k}^{l},r_{k}^{l},t_{k}^{l}),$ (6)
where $f(h,r,t)$ denotes the score function of the pretrained ComplEx model. $Sc(T_{k}^{l})$ measures the possibility of $T_{k}^{l}$ appearing in given KG. Further, we compute the average score of all triples occuring in the document $G_{k}$:
$Score(G_{k})=Avg(Sc(T_{k}^{l}),\ T_{k}^{l}\in\mathcal{T}_{k}).$ (7)
Document Selection.
Relying on the authenticity score of $G_{k}\in\mathcal{OI}$, we could finally obtain the selected documents. Specifically, considering the ranking design of Google search (Patil et al., 2021), we first choose the top-1 document $G_{1}$ in $\mathcal{OI}$. Then we select top n-1 documents from the remaining L-1 documents, through the authenticity score $Score(G_{k})$. Finally, we obtain the selected outside investigated information $\mathcal{OI^{\prime}}=\{G_{k},k=1,2,..,n\}$.